university-logo Introduction

Body Conclusion

Stuff About Bad Talks

Josh

May 1, 2008

university-logo Introduction

Body Conclusion

Outline

1 Introduction
2 Body

Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

3 Conclusion

university-logo Introduction

Body Conclusion

Why the Audience Should Care

I dunno. My professor told me to work on this problem. Here’s my first result: ….

university-logo Introduction

Body Conclusion

Let’s Try Again

Outlines are overrated.

Things you have to read have outlines. Things you want to read have plots.

Take the chance to grab your audience — give them a “plot synopsis”.

university-logo Introduction

Body Conclusion

Plot Synopsis

If you give a good talk, the audience will learn something without being bored or disctracted. There are some definite things you should avoid doing. Some of these apply particularly to math (and theoretical CS) talks.

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Outline

1 Introduction

2 Body

Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

3 Conclusion

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things to Avoid Under Any Circumstances While You Are Preparing to Speak

Dressing inappropriately

Minimum standards: clean, no holes or tears, no writing

Chewing gum (or anything else) Not knowing how to use your computer Not having your slides finished Not having your computer (or the room’s computer) ready when it’s time to go

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things to Avoid Under Any Circumstances While You Are Speaking

Random arm motions The presentation dance Reading the slide verbatim “Umm …. Oh sh*%& ….” Overly informal language Jargon

Technical talk is good, jargon is bad.

Talking too fast

Time your presentation beforehand. Adrenaline tends to make you speed up.

Going over time

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things to Avoid Under Any Circumstances In Your Slides

Titles which aren’t informative Titles which are too long Slides which don’t impart any information

Too many bells and whistles Excessive use of a laser pointer or mouse Typoes

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things to Avoid Under Any Circumstances In Your Slides

Titles which aren’t informative Titles which are too long Slides which don’t impart any information

Too many bells and whistles Excessive use of a laser pointer or mouse Typoes

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things to Avoid Under Any Circumstances In Your Slides

Titles which aren’t informative Titles which are too long Slides which don’t impart any information Too many bells and whistles

Excessive use of a laser pointer or mouse Typoes

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things to Avoid Under Any Circumstances In Your Slides

Titles which aren’t informative Titles which are too long Slides which don’t impart any information Too many bells and whistles

Excessive use of a laser pointer or mouse Typoes

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things to Avoid Under Any Circumstances In Your Slides

Titles which aren’t informative Titles which are too long Slides which don’t impart any information

Too many bells and whistles Excessive use of a laser pointer or mouse Typoes

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

More Things to Avoid Under Any Circumstances In Your Slides

Unreadable color combinations Changing colors for no reason

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things to Avoid in Any Kind of Twenty-Minute Talk

Too much detail

“The purpose of a 20-minute talk is to get the audience to read your paper.”

Too much time spent road-mapping Writing on the whiteboard/chalkboard

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Outline

1 Introduction

2 Body

Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

3 Conclusion

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Issues with Mathematical Content

Know your audience

Don’t underestimate, but… Don’t overestimate, either. Your audience will not be experts on your particular topic.

Some formulas are good, but not too many. Format your formulas correctly: E = mcˆ 2 vs. E = mc^{2}.

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Issues with Graphics

Graphics are good. Informative graphics are even better.

Include labels on graphs. For each graph, first say:

What the axes are. What this is a graph of.

Tables are not as good.

Tables (or text) that are too small are really bad. This often happens if you clip text out of another document.

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

POLICY FOR RESPONSIBLE USE OF ROSE-HULMAN COMPUTING FACILITIES
GENERAL GUIDELINES
This policy applies to all computing equipment, networks, network connections, software,
and services provided, maintained, or operated by the Waters Computing Center for Rose-Hulman.
In short, it applies to all Institute computing facilities and services, and to all users.
In some cases, it may be superseded or modified by policies of departments with their own
computing facilities. However, where departmental facilities are used for tasks that involve
facilities or services provided by the Waters Computing Center, this policy will apply.
Computing facilities are provided by the Institute to support and enhance the educational
experience for every student. They also are intended to assist the faculty's professional
activities and to provide administrative support for the Institute. Therefore, their misuse
harms every member of the Rose-Hulman community. Since everyone benefits from efficient
system management, the Computing Center staff encourages suggestions of ways to improve
operations.

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things that Mathematicians Tend to Dislike

Cutsey or gimmicky graphics, sound effects, animations Busy backgrounds Lots of overlays

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things that Mathematicians Tend to Dislike

Cutsey or gimmicky graphics, sound effects, animations Busy backgrounds Lots of overlays

university-logo

Introduction Body Conclusion  Things to Avoid Under Any Circumstances Issues with Math Talks in Particular

Things that Mathematicians Tend to Dislike

Cutsey or gimmicky graphics, sound effects, animations Busy backgrounds Lots of overlays

university-logo Introduction

Body Conclusion

What the Audience Should Take Away

Don’t repeat your introduction. Really don’t repeat your outline. What were the most important things you found out? What else would you like to do? How else could someone use the information?

university-logo Introduction

Body Conclusion

Question and Answer

If you don’t know the answer, don’t be afraid to say so. If you can answer partially, or a related question, go ahead. Refer to sources if appropriate. Don’t act like a question is stupid … But don’t act like you are too stupid for the question, either. Ask the questioner to rephrase if necessary.